How much money will there be in an account at the end of 5 years if $11000 is deposited at a 3.5% annual rate that is compounded continuously?
The amount after 5 years will be?

Answer by rfer(16322)

A=11000(2.7183)^0.035*5
A=11000(2.7183)^0.175
A=11000(1.1912476)
A=$13,103.72